Circular No. 5781

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993L IN IC 5270 N. Suntzeff, Cerro Tololo Interamerican Observatory, communicates:

“CCD images taken with the CTIO 0.9-m telescope by G. Williger on May 1.43 UT show SN 1993L to be 25′′ west and 1′′.4 north of the galaxy’s center. Reduction of the CCD frames by R. Aviles yields the following magnitudes (R and I on the Kron-Cousins system): B = 14.26, V = 13.60, R_{c} = 13.61, I_{c} = 13.71). The photometry was performed with a 14′′-diameter aperture and may be slightly contaminated by the galaxian light at the 0.03-magnitude level. The suspected supernova is in the disk of IC 5270 near a prominent dust lane, and the red colors may indicate high reddening. IC 5270 is a particularly interesting galaxy because it is a member of the Grus cluster and has been used in Tully-Fisher distance scale work, and further observations are urged.”

GRO J0422+32 B. A. Harmon and G. J. Fishman, Marshall Space Flight Center,

NASA; and W. S. Paciesas and M. S. Briggs, University of Alabama, Huntsville, report for the BATSE/CGRO Experiment Team: “Hard x-ray monitoring of GRO J0422+32 (cf. IAUC 5580, 5584, 5587, 5592, 5685, etc.) using Earth occultation indicated that flux from the object in the energy band 20-300 keV had decreased below the BATSE one-day detection limit of ∼ 100 mCrab (3σ) after 1993 Mar. 14. The characteristic times for exponential decay of the light curve (1/e) following the primary maximum (1992 mid-August; IAUC 5580, 5584) and the secondary maximum (1992 mid- to late-December; IAUC 5685) were similar (∼ 40 days). Spectral fits with an optically thin thermal bremsstrahlung model have shown that the spectrum hardened in the initial decay phase from kT ∼ 90 keV near the primary maximum in 1992 August to kT ∼ 115 keV around 1992 mid-September. No significant spectral changes were observed during the subsequent decline through 1993 March, which includes the secondary intensity maximum. Monitoring of GRO J0422+32 is continuing.”

V351 PUPPIS (NOVA PUPPIS 1991) Photometry by A. C. Gilmore, Mt. John University Observatory: Apr.

18.45 UT, V = 13.07 ± 0.02, U–B = –1.53 ± 0.04, B–V = +0.24 ± 0.00, V–R = +0.14 ± 0.03.

1993 May 1 Daniel W. E. Green